OpenStudy (anonymous):

equation of the line perpendicular to the line x-3y=9 and passing through the point (3,5)

OpenStudy (anonymous):

start converting your eq.in y=mx+c form . i.e. (x/3)-3=y ; here m=(1/3); m= slope of your line . you asked for perpendicular to this line . if two lines are perpendicular and their sopes are m and n thn ; m*n= (-1); so n=(-1/m); in your qu. n= -3; now you know slope and point . can find the eq easily.